OpenStudy (anonymous):

solve each equation for x 3^x =3^4 help

OpenStudy (mathmate):

\(\large 3^x=3^4\) Noting that \(\large log_3(3^x)=x\) and \(log_3(3^4)=4\) Take log to the base 3 on each side: \(\huge x=4\)

OpenStudy (anonymous):

Good job citizen. @mathmate

OpenStudy (mathmate):

Thank you! :)

OpenStudy (anonymous):

thanks how bout this and i will give u a medal 5^x-2=25

OpenStudy (mathmate):

@thegreatone215 I suggest you review the rules of priorities of operations, commonly known as PEDMAS, or BEMDAS or some other variants. Namely the order is parentheses, exponential, multiplication/division, THEN addition/subtraction. What you wrote would mean \(5^2-2=25\) which give answer as \(3\sqrt3\) But I suppose you actually mean: \(5^{x-2}=25\) which reduces to \(5^{x-2}=5^2\) and you can solve this equation in a way similar to the first one.

OpenStudy (anonymous):

yea i ment the second one on the computer it wont let me sqaure